graph each equation.
x2+y2-6x-4y+9=0

Answer by richwmiller(17219)

This is a circle.
x2+y2-6x-4y+9=0
x^2-6x+a+y^2-4y+b=-9
we need to complete the square twice to get the radius and center
6/2=3 3^2=9 add 9
4/2=2 2^4=4 add 4
(x-3)^2+(y-2)^2=4
center at (3,2) radius 2
some integer points for ease of plotting.
1,2
3,0
3,4
5,2